NoFly Corporation sells three different models of a mosquito “zapper.” Model A12 sells for $61 and has variable costs of $37. Model B22 sells for $110 and has variable costs of $78. Model C124 sells for $421 and has variable costs of $319. The sales mix of the three models is A12, 56%; B22, 29%; and C124, 15%. If the company has fixed costs of $274,504, how many units of each model must the company sell in order to break even? Also, what is the total break even point in units?
Answer #1

Solution:

Computation of weighted average contribution margin per unit
Particulars Model A12 Model B22 Model C124 Total
Selling price per unit $61.00 $110.00 $421.00
Variable cost per unit $37.00 $78.00 $319.00
Contribution margin per unit $24.00 $32.00 $102.00
Sales mix 56% 29% 15%
Weighted average contribution margin per unit $13.44 $9.28 $15.30 $38.02

Total breakeven units = Fixed costs / Weighted average contribution margin per unit = $274,504 / $38.02 = 7220 units

Breakeven units - Model A12 = 7220*56% = 4043 units

Breakeven units - Model B22 = 7220*29% = 2094 units

Breakeven units - C 124 = 7220*15% = 1083 units